OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x)) ? 2x – 2 3x + 2 4x + 2 2x + 6

OpenStudy (anonymous):

Just replace x in f(x) with 3x + 4. Mean, you have f(x) = x - 2 Here, in place of x put 3x + 4 and check what you will get.. Can you do it???

OpenStudy (foolaroundmath):

g(x) = 3x+4 so, f(g(x)) is essentially f(3x+4).. can you do this ?

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

Where you are facing problem in which step??

OpenStudy (anonymous):

seting it up

OpenStudy (anonymous):

See, what is g(x) given??

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

Guessing or you have solved it on your own??

OpenStudy (anonymous):

guessing

OpenStudy (anonymous):

Why?? You don't want to solve this to get the answer???

OpenStudy (anonymous):

lol oh sorry

OpenStudy (anonymous):

Then??

OpenStudy (anonymous):

can you help me set it up

OpenStudy (anonymous):

Oh sure why not.. Keep on answering what I am asking you will definitely set it up... What is value of g(x) given in problem??

OpenStudy (anonymous):

3x+4

OpenStudy (anonymous):

Yes and what is f(x) given??

OpenStudy (anonymous):

x-2

OpenStudy (anonymous):

And what we are to be found??

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

-2(3x+4)

OpenStudy (anonymous):

So, now can I write f(g(x)) as f(3x + 4)?? because you know what is the value of g(x).. Got or not??

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and what do we put in for f

OpenStudy (anonymous):

f(g(x)) = f o g(x) it means that g(x) is subtitute in x-> f(x).. understand?

OpenStudy (anonymous):

f(x) = x - 2 f(3x + 4) = ???? We will put nothing in place of f, it is just representing that it is a function..named as f.. In f(x), In place of x if I put 3x + 4, then what will you get??

OpenStudy (anonymous):

which x?

OpenStudy (anonymous):

In this "f(x)". you are seeing x??

OpenStudy (anonymous):

im lost..

OpenStudy (anonymous):

I make you see by a drawing...

OpenStudy (anonymous):